# Learning Patient-Specific Disease Dynamics with Latent Flow Matching for Longitudinal Imaging Generation

### v1

@@ 1 Introduction
Modeling disease progression is a fundamental problem in healthcare (Suk & Shen, 2013; Wang et al., 2020), with profound implications for early detection and individualized treatment. Disease does not evolve in a uniform manner: differences in anatomy, genetics, and environment give rise to heterogeneous trajectories, which are an intrinsic property of disease dynamics (Real & Biek, 2007; Parratt et al., 2016; Li et al., 2025). Accounting for this heterogeneity is essential for accurately characterizing temporal changes, identifying preclinical biomarkers, and designing personalized therapies that can improve clinical outcomes (Frisoni et al., 2010). However, most progression models capture only population-level trends, overlooking individualized variation (Rokuss et al., 2025; Yang et al., 2025b). This limitation obscures early-stage signals and misaligns disease severity with underlying anatomy, ultimately reducing clinical utility. Capturing patient-specific dynamics is therefore critical, but remains an underexplored challenge (Lai et al., 2025).
Recent advances in generative modeling have opened new opportunities for studying disease progression. Traditional biomarker-based approaches often reduce complexity to coarse measures (e.g.,  volume changes (Jack Jr et al., 2003; Tabrizi et al., 2011) or disease rating (O’Bryant et al., 2008; Young et al., 2020)), which can obscure heterogeneous disease dynamics. In contrast, longitudinal medical image generation offers a principled framework for visualizing symptomatic transitions and individualized patterns of anatomical change. Rather than limiting analysis to numeric outcomes, synthesized images yield interpretable and more informative visual representations that support structured analyses and offer more actionable insights for clinical practice (Chen et al., 2024a; Wu et al., 2024; Chen et al., 2024b).
Several methods have attempted to model disease progression. For instance, DiffuseMorph (Kim et al., 2022) introduces a morph field to describe pixel-wise displacements, and TADM (Litrico et al., 2024) incorporates temporal awareness by predicting residual images conditioned on age gaps. Despite these advances, generating temporally consistent and individualized disease trajectories remains an open problem. Beyond high-fidelity image synthesis, it is crucial for generative models to preserve fine-grained anatomical structures and accurately capture patient-specific dynamics. Although BrLP (Puglisi et al., 2024) takes a step toward personalization by modeling volume changes, such guidance is coarse and provides only limited control over individualized trajectories.
@@ 2 Related Work
Generative model for disease progression. Generative modeling has emerged as a powerful strategy for simulating longitudinal disease changes. Early efforts relied on generative adversarial networks (GANs), either by introducing morphological priors to simulate aging effects (Ravi et al., 2019) or by predicting deformation fields with 3D conditional GANs instead of direct intensity values (Ravi et al., 2022).
More recently, diffusion models (Ho et al., 2020) have become the dominant paradigm due to their superior fidelity and stability. Extensions of this framework introduce temporal awareness and anatomical constraints. Sequence-Aware Diffusion Model (SADM) leverages multiple scans to autoregressively generate future MRIs (Yoon et al., 2023), while DiffuseMorph (Kim et al., 2022) produces smooth voxel-wise deformation fields across timepoints, thereby preserving topology. Progressive Image Editing (PIE) (Liang et al., 2023) edits disease-related regions in medical images using text-guided diffusion models to reflect how a patient’s condition evolves over time. Kyung et al. (Kyung et al., 2024) adopt electronic health records (EHRs), including medical history, to model temporal changes and enable medically meaningful generation. Temporally-Aware Diffusion Model (TADM) conditions on age maps and adopt a residual prediction strategy, modeling progression as incremental changes rather than absolute reconstructions (Litrico et al., 2024).
Conditioning strategies further enhance subject specificity and anatomical fidelity. Brain Latent Progression (BrLP) employs ControlNet with volumetric ratio conditioning to synthesize individualized disease trajectories while integrating population priors (Puglisi et al., 2024; 2025). Similarly, BrainMRDiff fuses multiple structural masks (e.g., white matter, gray matter, ventricles) into a unified conditioning representation, substantially improving anatomical realism (Bhattacharya et al., 2025). Imageflownet (Liu et al., 2025a) learns multiscale joint patient embeddings and position-parameterized neural flow fields to forecast full image-level trajectories.
@@
Subject-level consistency and biomedical plausibility. Ensuring subject-specific consistency and biomedical plausibility is a central challenge in modeling synthetic disease progression. Early approaches such as the Longitudinal Variational Autoencoder (Ramchandran et al., 2021) and Longitudinal Self-Supervised Learning (LNE) (Zhao et al., 2021) focus on learning patient-temporal dynamics by constructing temporally coherent latent spaces, thereby producing individualized representations for downstream tasks.
Subsequent works introduced explicit mechanisms to regularize longitudinal consistency. Ren et al. (Ren et al., 2023) enforce subject-level coherence through local and multi-scale spatio-temporal features within a self-supervised framework. In a complementary direction, Treatment-Aware Diffusion (TaDiff) (Liu et al., 2025b) incorporates prior scans and treatment information to jointly generate future MRIs and tumor segmentations, thereby improving biological plausibility by conditioning on clinically relevant variables.
@@ 3.1 Preliminaries
Flow matching (Lipman et al., 2023; Tong et al., 2024) trains continuous normalizing flows by directly aligning vector fields instead of relying on stochastic denoising objectives as in diffusion models. Let $\pi_{0}$ and $\pi_{1}$ denote the source and target distributions, respectively. Flow matching defines a family of intermediate distributions $\{\pi_{t}\}_{t\in[0,1]}$ along a continuous trajectory connecting $\pi_{0}$ and $\pi_{1}$. The goal is to learn a time-dependent velocity field $v_{\theta}(\mathbf{x},t)$ such that the solution of the ordinary differential equation (ODE) (Lu et al., 2022) transports $\pi_{0}$ exactly to $\pi_{1}$ at $t=1$:
@@
Unlike score-based diffusion models, which approximate the score function $\nabla_{\mathbf{x}}\log\pi_{t}(\mathbf{x})$ via noisy denoising objectives, flow matching bypasses score estimation by directly regressing the neural vector field $v_{\theta}(\mathbf{x},t)$ onto analytically known target fields. For example, under conditional flow matching (Lipman et al., 2023), one defines a reference path $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+t\mathbf{x}_{1}$ between samples $\mathbf{x}_{0}\sim\pi_{0}$ and $\mathbf{x}_{1}\sim\pi_{1}$, yielding a target velocity
@@
To further incorporate stochasticity (Albergo et al., 2023), the deterministic flow can be extended to a stochastic differential equation of the form
@@ 3.2 ArcRank: Longitudinal Alignment of Patient Latent
To enhance sensitivity to patient-specific dynamics, we introduce a chronologically-aware latent representation learned via a new contrastive objective, termed ArcRank Loss. This objective aligns representations across time for the same individual by enforcing both angular consistency and temporal ordering in the embedding space. Let $\mathbf{z}$ denote the latent representation of an input $\mathbf{x}$, extracted from a Variational Autoencoder (VAE) (Kingma & Welling, 2013) encoder $E(\cdot)$. The loss encourages
@@
In practice, we extract these components via Singular Value Decomposition (SVD) (Klema & Laub, 1980) and write
@@ 4.1 Experiment Settings
We build our study on three major Alzheimer’s disease (AD) cohorts: ADNI (Petersen et al., 2010), OASIS-3 (LaMontagne et al., 2019), and AIBL (Ellis et al., 2009). Every MRI undergoes the same preprocessing pipeline: N4 bias correction (Tustison et al., 2010), skull stripping (Hoopes et al., 2022), resampling to 1.5 mm3 voxels. We then normalize tissue contrasts and register all scans to a refined standard brain template. We randomly divide data into 80% training, 5% validation, and 15% test.
We benchmark against a broad spectrum of recent approaches. Direct-prediction methods include CardiacAging (Campello et al., 2022) and CounterSynth (Pombo et al., 2023). Deformation-based modeling represented by DiffuseMorph (Kim et al., 2022). Diffusion-based methods span multiple variants: SADM (Yoon et al., 2023), which denoises full images; TADM (Litrico et al., 2024), which denoises residuals; and two ControlNet-based (Zhang et al., 2023) models, BrLP (Puglisi et al., 2024) and MambaControl (Yang et al., 2025a).
Metrics. We evaluate reconstruction quality with two standard image-level metrics: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM). To capture anatomical fidelity, we followed BrLP (Puglisi et al., 2024) to report region-level mean absolute error (MAE) over clinically relevant structures (hippocampus, amygdala, lateral ventricles, cerebrospinal fluid (CSF), and thalamus).
@@
CardiacAging (Campello et al., 2022) Direct 27.78 $\pm$ 1.49 92.04 $\pm$ 0.99 28.41 $\pm$ 1.41 90.30 $\pm$ 0.90 26.23 $\pm$ 1.45 86.17 $\pm$ 2.06
CounterSynth (Pombo et al., 2023) Direct 28.24 $\pm$ 1.31 92.84 $\pm$ 0.95 28.96 $\pm$ 1.25 92.10 $\pm$ 0.91 26.79 $\pm$ 1.42 86.45 $\pm$ 1.98
DiffuseMorph (Kim et al., 2022) Deform 29.56 $\pm$ 1.63 93.57 $\pm$ 0.93 29.17 $\pm$ 1.54 92.70 $\pm$ 1.02 28.13 $\pm$ 1.36 88.16 $\pm$ 1.69
SADM (Yoon et al., 2023) Noise 26.94 $\pm$ 2.28 85.15 $\pm$ 2.72 27.97 $\pm$ 1.91 89.40 $\pm$ 1.78 26.74 $\pm$ 1.80 86.36 $\pm$ 2.37
@@
ImageFlowNet Liu et al. (2025a) Flow 28.37 $\pm$ 1.18 92.96 $\pm$ 0.88 29.08 $\pm$ 1.12 92.23 $\pm$ 0.84 27.92 $\pm$ 1.29 87.63 $\pm$ 1.82
BrLP (Puglisi et al., 2024) Noise 28.51 $\pm$ 1.77 91.52 $\pm$ 1.31 28.96 $\pm$ 1.36 92.21 $\pm$ 0.97 27.98 $\pm$ 1.40 87.93 $\pm$ 1.44
MambaControl (Yang et al., 2025a) Noise 29.72 $\pm$ 1.04 93.60 $\pm$ 0.96 29.86 $\pm$ 1.21 93.17 $\pm$ 0.98 28.24 $\pm$ 1.28 88.30 $\pm$ 1.52
@@
CardiacAging (Campello et al., 2022) Direct 0.289 $\pm$ 0.33 0.771 $\pm$ 0.12 0.281 $\pm$ 0.35 0.698 $\pm$ 0.13 0.345 $\pm$ 0.35 0.739 $\pm$ 0.13
CounterSynth (Pombo et al., 2023) Direct 0.272 $\pm$ 0.32 0.704 $\pm$ 0.11 0.256 $\pm$ 0.30 0.672 $\pm$ 0.12 0.298 $\pm$ 0.34 0.690 $\pm$ 0.12
DiffuseMorph (Kim et al., 2022) Deform 0.230 $\pm$ 0.28 0.516 $\pm$ 0.10 0.226 $\pm$ 0.27 0.482 $\pm$ 0.10 0.279 $\pm$ 0.29 0.503 $\pm$ 0.11
SADM (Yoon et al., 2023) Noise 0.283 $\pm$ 0.34 0.746 $\pm$ 0.11 0.278 $\pm$ 0.30 0.693 $\pm$ 0.12 0.365 $\pm$ 0.39 0.728 $\pm$ 0.12
@@
ImageFlowNet Liu et al. (2025a) Flow 0.259 $\pm$ 0.30 0.589 $\pm$ 0.11 0.240 $\pm$ 0.29 0.561 $\pm$ 0.12 0.283 $\pm$ 0.32 0.574 $\pm$ 0.12
BrLP (Puglisi et al., 2024) Noise 0.257 $\pm$ 0.32 0.630 $\pm$ 0.10 0.278 $\pm$ 0.31 0.594 $\pm$ 0.11 0.335 $\pm$ 0.36 0.622 $\pm$ 0.11
MambaControl (Yang et al., 2025a) Noise 0.225 $\pm$ 0.30 0.554 $\pm$ 0.09 0.249 $\pm$ 0.30 0.525 $\pm$ 0.10 0.299 $\pm$ 0.32 0.561 $\pm$ 0.10
@@ 4.3 Ablation study
The baseline Latent Flow Matching (LFM) (Dao et al., 2023) model without conditioning performs the worst across all metrics, showing notably low SSIM/PSNR and high error values. Incorporating conditional information (Conditional LFM) significantly improves image quality (SSIM from 88.62 to 91.20; PSNR from 27.59 to 28.46) and reduces both Region MAE and $\Delta$-RMAE.

### v2

@@ 1 Introduction
Modeling disease progression is a fundamental problem in healthcare (suk2013deep; wang2020deep), with profound implications for early detection and individualized treatment. Disease does not evolve in a uniform manner: differences in anatomy, genetics, and environment give rise to heterogeneous trajectories, which are an intrinsic property of disease dynamics (real2007spatial; parratt2016infectious; li2025pants). Accounting for this heterogeneity is essential for accurately characterizing temporal changes, identifying preclinical biomarkers, and designing personalized therapies that can improve clinical outcomes (frisoni2010clinical). However, most progression models capture only population-level trends, overlooking individualized variation (rokuss2025lesionlocator; yang2025medical). This limitation obscures early-stage signals and misaligns disease severity with underlying anatomy, ultimately reducing clinical utility. Capturing patient-specific dynamics is therefore critical, but remains an underexplored challenge (lai2025patient).
Recent advances in generative modeling have opened new opportunities for studying disease progression. Traditional biomarker-based approaches often reduce complexity to coarse measures (e.g.,  volume changes (jack2003mri; tabrizi2011biological) or disease rating (o2008staging; young2020imaging)), which can obscure heterogeneous disease dynamics. In contrast, longitudinal medical image generation offers a principled framework for visualizing symptomatic transitions and individualized patterns of anatomical change. Rather than limiting analysis to numeric outcomes, synthesized images yield interpretable and more informative visual representations that support structured analyses and offer more actionable insights for clinical practice (chen2024towards; wu2024freetumor; chen2024analyzing).
Several methods have attempted to model disease progression. For instance, DiffuseMorph (kim2022diffusemorph) introduces a morph field to describe pixel-wise displacements, and TADM (litrico2024tadm) incorporates temporal awareness by predicting residual images conditioned on age gaps. Despite these advances, generating temporally consistent and individualized disease trajectories remains an open problem. Beyond high-fidelity image synthesis, it is crucial for generative models to preserve fine-grained anatomical structures and accurately capture patient-specific dynamics. Although BrLP (puglisi2024enhancing) takes a step toward personalization by modeling volume changes, such guidance is coarse and provides only limited control over individualized trajectories.
@@ 2 Related Work
Generative model for disease progression. Generative modeling has emerged as a powerful strategy for simulating longitudinal disease changes. Early efforts relied on generative adversarial networks (GANs), either by introducing morphological priors to simulate aging effects (raviDegenerativeAdversarialNeuroImage2019a) or by predicting deformation fields with 3D conditional GANs instead of direct intensity values (ravi2022degenerative).
More recently, diffusion models (ddpm) have become the dominant paradigm due to their superior fidelity and stability. Extensions of this framework introduce temporal awareness and anatomical constraints. Sequence-Aware Diffusion Model (SADM) leverages multiple scans to autoregressively generate future MRIs (yoon2023sadm), while DiffuseMorph (kim2022diffusemorph) produces smooth voxel-wise deformation fields across timepoints, thereby preserving topology. Progressive Image Editing (PIE) (liang2023pie) edits disease-related regions in medical images using text-guided diffusion models to reflect how a patient’s condition evolves over time. Kyung et al. (kyung2024towards) adopt electronic health records (EHRs), including medical history, to model temporal changes and enable medically meaningful generation. Temporally-Aware Diffusion Model (TADM) conditions on age maps and adopt a residual prediction strategy, modeling progression as incremental changes rather than absolute reconstructions (litrico2024tadm).
Conditioning strategies further enhance subject specificity and anatomical fidelity. Brain Latent Progression (BrLP) employs ControlNet with volumetric ratio conditioning to synthesize individualized disease trajectories while integrating population priors (puglisi2024enhancing; puglisi2025brain). Similarly, BrainMRDiff fuses multiple structural masks (e.g., white matter, gray matter, ventricles) into a unified conditioning representation, substantially improving anatomical realism (bhattacharya2025brainmrdiff). Imageflownet (liu2025imageflownet) learns multiscale joint patient embeddings and position-parameterized neural flow fields to forecast full image-level trajectories.
@@
Subject-level consistency and biomedical plausibility. Ensuring subject-specific consistency and biomedical plausibility is a central challenge in modeling synthetic disease progression. Early approaches such as the Longitudinal Variational Autoencoder (ramchandran2021longitudinalvariationalautoencoder) and Longitudinal Self-Supervised Learning (LNE) (zhao2021longitudinal) focus on learning patient-temporal dynamics by constructing temporally coherent latent spaces, thereby producing individualized representations for downstream tasks.
Subsequent works introduced explicit mechanisms to regularize longitudinal consistency. Ren et al. (ren2023localspatiotemporalrepresentationlearning) enforce subject-level coherence through local and multi-scale spatio-temporal features within a self-supervised framework. In a complementary direction, Treatment-Aware Diffusion (TaDiff) (liu2025treatment) incorporates prior scans and treatment information to jointly generate future MRIs and tumor segmentations, thereby improving biological plausibility by conditioning on clinically relevant variables.
@@ 3.1 Preliminaries
Flow matching (lipman2023flow; tong2023improving) trains continuous normalizing flows by directly aligning vector fields instead of relying on stochastic denoising objectives as in diffusion models. Let $\pi_{0}$ and $\pi_{1}$ denote the source and target distributions, respectively. Flow matching defines a family of intermediate distributions $\{\pi_{t}\}_{t\in[0,1]}$ along a continuous trajectory connecting $\pi_{0}$ and $\pi_{1}$. The goal is to learn a time-dependent velocity field $v_{\theta}(\mathbf{x},t)$ such that the solution of the ordinary differential equation (ODE) (lu2022dpm) transports $\pi_{0}$ exactly to $\pi_{1}$ at $t=1$:
@@
Unlike score-based diffusion models, which approximate the score function $\nabla_{\mathbf{x}}\log\pi_{t}(\mathbf{x})$ via noisy denoising objectives, flow matching bypasses score estimation by directly regressing the neural vector field $v_{\theta}(\mathbf{x},t)$ onto analytically known target fields. For example, under conditional flow matching (lipman2023flow), one defines a reference path $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+t\mathbf{x}_{1}$ between samples $\mathbf{x}_{0}\sim\pi_{0}$ and $\mathbf{x}_{1}\sim\pi_{1}$, yielding a target velocity
@@
To further incorporate stochasticity (albergo2023stochastic), the deterministic flow can be extended to a stochastic differential equation of the form
@@ 3.2 ArcRank: Longitudinal Alignment of Patient Latent
To enhance sensitivity to patient-specific dynamics, we introduce a chronologically-aware latent representation learned via a new contrastive objective, termed ArcRank Loss. This objective aligns representations across time for the same individual by enforcing both angular consistency and temporal ordering in the embedding space. Let $\mathbf{z}$ denote the latent representation of an input $\mathbf{x}$, extracted from a Variational Autoencoder (VAE) (kingma2013auto) encoder $E(\cdot)$. The loss encourages
@@
In practice, we extract these components via Singular Value Decomposition (SVD) (klema1980singular) and write
@@ 4.1 Experiment Settings
We build our study on three major Alzheimer’s disease (AD) cohorts: ADNI (petersen2010alzheimer), OASIS-3 (lamontagne2019oasis), and AIBL (ellis2009australian). Every MRI undergoes the same preprocessing pipeline: N4 bias correction (tustison2010n4itk), skull stripping (hoopes2022synthstrip), resampling to 1.5 mm3 voxels. We then normalize tissue contrasts and register all scans to a refined standard brain template. We randomly divide data into 80% training, 5% validation, and 15% test.
We benchmark against a broad spectrum of recent approaches. Direct-prediction methods include CardiacAging (campello2022cardiac) and CounterSynth (pombo2023equitable). Deformation-based modeling represented by DiffuseMorph (kim2022diffusemorph). Diffusion-based methods span multiple variants: SADM (yoon2023sadm), which denoises full images; TADM (litrico2024tadm), which denoises residuals; and two ControlNet-based (zhang2023adding) models, BrLP (puglisi2024enhancing) and MambaControl (yang2025mambacontrol).
Metrics. We evaluate reconstruction quality with two standard image-level metrics: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM). To capture anatomical fidelity, we followed BrLP (puglisi2024enhancing) to report region-level mean absolute error (MAE) over clinically relevant structures (hippocampus, amygdala, lateral ventricles, cerebrospinal fluid (CSF), and thalamus).
@@
CardiacAging (campello2022cardiac) Direct 27.78 $\pm$ 1.49 92.04 $\pm$ 0.99 28.41 $\pm$ 1.41 90.30 $\pm$ 0.90 26.23 $\pm$ 1.45 86.17 $\pm$ 2.06
CounterSynth (pombo2023equitable) Direct 28.24 $\pm$ 1.31 92.84 $\pm$ 0.95 28.96 $\pm$ 1.25 92.10 $\pm$ 0.91 26.79 $\pm$ 1.42 86.45 $\pm$ 1.98
DiffuseMorph (kim2022diffusemorph) Deform 29.56 $\pm$ 1.63 93.57 $\pm$ 0.93 29.17 $\pm$ 1.54 92.70 $\pm$ 1.02 28.13 $\pm$ 1.36 88.16 $\pm$ 1.69
SADM (yoon2023sadm) Noise 26.94 $\pm$ 2.28 85.15 $\pm$ 2.72 27.97 $\pm$ 1.91 89.40 $\pm$ 1.78 26.74 $\pm$ 1.80 86.36 $\pm$ 2.37
@@
ImageFlowNet liu2025imageflownet Flow 28.37 $\pm$ 1.18 92.96 $\pm$ 0.88 29.08 $\pm$ 1.12 92.23 $\pm$ 0.84 27.92 $\pm$ 1.29 87.63 $\pm$ 1.82
BrLP (puglisi2024enhancing) Noise 28.51 $\pm$ 1.77 91.52 $\pm$ 1.31 28.96 $\pm$ 1.36 92.21 $\pm$ 0.97 27.98 $\pm$ 1.40 87.93 $\pm$ 1.44
MambaControl (yang2025mambacontrol) Noise 29.72 $\pm$ 1.04 93.60 $\pm$ 0.96 29.86 $\pm$ 1.21 93.17 $\pm$ 0.98 28.24 $\pm$ 1.28 88.30 $\pm$ 1.52
@@
CardiacAging (campello2022cardiac) Direct 0.289 $\pm$ 0.33 0.771 $\pm$ 0.12 0.281 $\pm$ 0.35 0.698 $\pm$ 0.13 0.345 $\pm$ 0.35 0.739 $\pm$ 0.13
CounterSynth (pombo2023equitable) Direct 0.272 $\pm$ 0.32 0.704 $\pm$ 0.11 0.256 $\pm$ 0.30 0.672 $\pm$ 0.12 0.298 $\pm$ 0.34 0.690 $\pm$ 0.12
DiffuseMorph (kim2022diffusemorph) Deform 0.230 $\pm$ 0.28 0.516 $\pm$ 0.10 0.226 $\pm$ 0.27 0.482 $\pm$ 0.10 0.279 $\pm$ 0.29 0.503 $\pm$ 0.11
SADM (yoon2023sadm) Noise 0.283 $\pm$ 0.34 0.746 $\pm$ 0.11 0.278 $\pm$ 0.30 0.693 $\pm$ 0.12 0.365 $\pm$ 0.39 0.728 $\pm$ 0.12
@@
ImageFlowNet liu2025imageflownet Flow 0.259 $\pm$ 0.30 0.589 $\pm$ 0.11 0.240 $\pm$ 0.29 0.561 $\pm$ 0.12 0.283 $\pm$ 0.32 0.574 $\pm$ 0.12
BrLP (puglisi2024enhancing) Noise 0.257 $\pm$ 0.32 0.630 $\pm$ 0.10 0.278 $\pm$ 0.31 0.594 $\pm$ 0.11 0.335 $\pm$ 0.36 0.622 $\pm$ 0.11
MambaControl (yang2025mambacontrol) Noise 0.225 $\pm$ 0.30 0.554 $\pm$ 0.09 0.249 $\pm$ 0.30 0.525 $\pm$ 0.10 0.299 $\pm$ 0.32 0.561 $\pm$ 0.10
@@ 4.3 Ablation study
The baseline Latent Flow Matching (LFM) (dao2023flow) model without conditioning performs the worst across all metrics, showing notably low SSIM/PSNR and high error values. Incorporating conditional information (Conditional LFM) significantly improves image quality (SSIM from 88.62 to 91.20; PSNR from 27.59 to 28.46) and reduces both Region MAE and $\Delta$-RMAE.